OpenStudy (anonymous):

Solve this equation? e^(2^x) = (e^3x) (1/(e^2))

OpenStudy (anonymous):

\[e^{2^x}=\frac{e^{3x}}{e^2}\]?

OpenStudy (anonymous):

if so, right hand side is \[e^{3x-2}\] and since exp is a one to one function your job is to solve \[2^x=3x-2\]

OpenStudy (anonymous):

unless the left hand hand side is \[(e^2)^x\]

OpenStudy (anonymous):

in which case you solve \[2x=3x-2\]

OpenStudy (anonymous):

in which case you get x = 2 by inspection

OpenStudy (anonymous):

actually 2 is a solution to \[2^x=3x-2\] as well, but there is also another solution that i have no idea how to find using algebra

OpenStudy (anonymous):

yes the first one is how it looks

OpenStudy (anonymous):

hello sensei! did you see that my myininaya is a \[\color{red}{\text{master!}}\]

OpenStudy (anonymous):

well if it is the first one 2 is a solution. there is another solution that i can find graphically. maybe there is some other way to find it. that is there are two solutions to \[2^x=3x-2\]

OpenStudy (amistre64):

i saw :) they grow up so fast these days

myininaya (myininaya):

lol

OpenStudy (anonymous):

we are all following in your footsteps

OpenStudy (amistre64):

i hope not lol

OpenStudy (anonymous):

hey how about \[\color{green}{\text{children of amistre}}\] as the next level?

OpenStudy (amistre64):

with any luck yall do better :)

myininaya (myininaya):

amistre how did you become moderator? and do you have any cool chores?

myininaya (myininaya):

like do they make you get the e-pizza?

OpenStudy (amistre64):

its by invitations; not really an experience level. my chores are to pretty much try to encourage safe interneting .. and take out the trash

myininaya (myininaya):

so you don't get the e-pizza? thats lame

OpenStudy (amistre64):

they make me grovel and beg for a window seat :)

myininaya (myininaya):

lol